# NeRFmentation: NeRF-based Augmentation for Monocular Depth Estimation

### v1

@@ Abstract
The capabilities of monocular depth estimation (MDE) models are limited by the availability of sufficient and diverse datasets. In the case of MDE models for autonomous driving, this issue is exacerbated by the linearity of the captured data trajectories. We propose a NeRF-based data augmentation pipeline to introduce synthetic data with more diverse viewing directions into training datasets and demonstrate the benefits of our approach to model performance and robustness. Our data augmentation pipeline, which we call “NeRFmentation”, trains NeRFs on each scene in the dataset, filters out subpar NeRFs based on relevant metrics, and uses them to generate synthetic RGB-D images captured from new viewing directions. In this work, we apply our technique in conjunction with three state-of-the-art MDE architectures on the popular autonomous driving dataset, KITTI, augmenting its training set of the Eigen split. We evaluate the resulting performance gain on the original test set, a separate popular driving set, and our own synthetic test set.
@@ 1 Introduction
Figure 1: NeRFmentation. We develop a novel dataset augmentation pipeline that utilizes NeRF-generated data rendered from poses unseen in the original dataset to improve the robustness of Monocular Depth Estimation networks. The method is intended to improve generalization capability in cases where the initial dataset provides limited spatial variability. The RGB image is taken from Waymo Open Dataset [35]
@@
As the field of computer vision continues to advance at a rapid pace, the pursuit of safer and more reliable autonomous driving systems remains paramount. Within this context, monocular depth estimation (MDE) plays an important role, offering a key solution to the complex challenge of depth perception in dynamic road environments. Although it offers a solution to these particular situations, it comes with extra challenges compared to stereo or in general multi-camera depth estimation. Accurately estimating the depth from a single image remains a challenging problem due to the inherent scale ambiguity in 2D projections of 3D scenes, which leads to an infinite number of geometrically plausible 3D reprojections for every 2D image. This means that monocular depth estimation is an underdetermined problem that relies on extracting and interpreting visual cues in images correctly. To address this challenge, numerous approaches leverage large-scale datasets to train deep neural networks, allowing them to learn to interpret visual cues and transform them into complex mappings between image features and corresponding depth information.
Figure 2: Overview of our proposed pipeline. The pipeline consists of four distinct stages. We train a NeRF for each scene in the MDE training dataset, reserving a portion of the images and poses per scene to evaluate NeRF reconstruction quality. Subpar scene NeRFs are filtered out. Then, we render novel views of the scene by systematically perturbing the original poses and querying the NeRF with these poses. Lastly, we combine the novel views with the original views of the scene to create a NeRFmented training dataset for the MDE network. Here, the source dataset is KITTI. [12]
This mapping is implicitly learned from the data that is being fed into the system during training time and a common approach to making the model learn these mappings is to implement sophisticated modules to extend the underlying neural network architecture.
While improving the model architecture is a promising approach to this problem, it also introduces the need for more training data due to an increased model complexity. This need for more training data can be dealt with by adding an existing dataset or by recording a new one. The former is common practice [22], although there are only finitely many real-world datasets. Recording a new dataset, however, is in the majority of cases infeasible since this introduces massive costs. This infeasibility of recording new datasets and the limited availability of real-world datasets leads to the need for different data sources. For that, one can resort to using synthetic datasets [30, 21, 23, 26]. This could lead to an unwanted domain shift which in turn can be compensated by using techniques from domain adaptation [39, 11, 25] to modify the data to match the training distribution. Domain adaptation by itself brings its own limitations as there is no guarantee that it matches the target distribution.
@@
In recent years, however, Neural Radiance Field (NeRF) methods that can encode complete 3D scenes have emerged. These NeRFs have the ability to generate photorealistic and geometrically consistent, dense RGB-D images from novel viewpoints with remarkable accuracy.
We propose to use the capabilities of NeRFs to generate high-quality RGB-D images from novel viewpoints in order to enlarge and diversify existing real-world datasets. This method is able to match the appearance distribution of the source dataset while generating novel viewpoints that would be arduous and costly to capture by recording a new dataset. Thus, we alleviate the need for combining synthetic datasets and domain adaptation as well as making MDE models more robust to images taken from out-of-distribution poses.
These benefits come with the extra advantage that the rendering of novel views can be automated and is very fast, while only relying upon reasonable computing power which is needed to train MDE networks anyway.
Moreover, this approach does not require designing a new model architecture, as we show that our method makes existing MDE network architectures more robust while keeping comparable performance or improving slightly with respect to the baseline.
Our main contributions are as follows:
  • We propose a novel dataset augmentation scheme for Monocular Depth Estimation that leverages NeRFs to augment a limited-scale dataset via scene reconstruction and virtual view RGB-D image generation for supervised training of data-driven MDE networks.
  • We conduct extensive experiments on multiple outdoor datasets, showing its robust performance gains on the target dataset, while keeping comparable performance on the source dataset which validates the effectiveness of our approach, showcasing the benefits of our scheme to MDE networks.
@@ 2 Related work
Figure 3: Qualitative comparison of Depth-Nerfacto vs. Depth-Nerfacto-Huge reconstruction on KITTI [12]. The figure shows reconstructions of training images using both Depth-Nerfacto and Depth-Nerfacto-Huge. Depth-Nerfacto-Huge outputs exhibit higher levels of detail and better accuracy in both the generated RGB and depth images compared to Depth-Nerfacto.
Monocular Depth Estimation.
The task of Monocular Depth Estimation (MDE) involves predicting the depth values for each pixel given a monocular 2D image. Due to the inherent difficulty in predicting the correct 3D reprojection to a 2D image, well-performing models employ sophisticated data-driven methods to infer the depth values. Initially, CNN-based encoder-decoder architectures [24, 9] that interpreted the MDE task as a pure regression task were employed. With the introduction of transformers [38] and their subsequent introduction into Computer Vision as vision transformers (ViT) [8], methods that employed a combination of CNNs and ViTs emerged, such as [6, 15, 16]. AdaBins [6] is a popular example of this class of models which interprets MDE as an ordinal regression problem similar to [10] but with the addition of learning adaptive bin centers and sizes. For this, it includes a ViT [8] into the CNN itself. In contrast to AdaBins [6], DepthFormer [16] treats MDE as a pure regression problem and employs the ViT [8] to directly encode images, purportedly exploiting the ViT’s strengths. Additionally, Li et al. introduce the hierarchical aggregation and heterogeneous interaction (HAHI) module, to enhance ViT features and leverage the affinity between ViT and convolution features [16]. BinsFormer [15] follows the idea of combining ordinal regression on binned depth and transformer networks. We do not develop a novel MDE architecture, but instead, use AdaBins [6], DepthFormer [16], and BinsFormer [15] to show the transferability and versatility of our method.
Neural Radiance Fields (NeRFs).
NeRFs are a scene representation method introduced in the seminal work [18] by Mildenhall et al. Using a sparse set of input views, they optimize an underlying volumetric scene function, taking advantage of a differentiable rendering pipeline. Novel views as well as depth data can then be synthesized by querying the network with novel camera poses. Nerfstudio [37] is a framework that facilitates the design, training, and testing of a wide variety of NeRF-based architectures through a modular approach. Nerfacto is an extension of the original NeRF developed by the Nerfstudio team. This model leverages the modularity of Nerfstudio to introduce new features proposed in recent NeRF literature such as camera pose refinement, per-image appearance conditioning, proposal sampling, scene contraction, and hash encoding. This combination of features leads Nerfacto to achieve near-SOTA novel view quality at a significantly faster speed than competitors such as MipNeRF [3]. Among further extensions of Nerfacto is depth-Nerfacto, extending this model with depth supervision, by providing a new loss term as proposed by Deng et al. [7]. Finally, Nerfacto-huge is the largest in a series of Nerfacto models offering different variants with different model sizes and capacities, offering the highest fidelity scene representations among all Nerfacto variants. When choosing an appropriate NeRF variant in our work, we once again leverage the modularity and customizability exposed by Nerfstudio to combine the aforementioned extensions of Nerfacto to benefit from both their advantages.
In the spirit of using NeRFs not only for novel view synthesis of RGBs but also other properties, MuVieNeRF [42] extends Neural Radiance Fields to surface normal prediction, shading estimation, edges detection, keypoints detection, and semantic labeling. The authors argue that exploiting the pose data for predictions leads to better results than when using traditional CNNs.

### v2

# NeRFmentation: Improving Monocular Depth Estimation with NeRF-based Data Augmentation

@@ Abstract
The capabilities of monocular depth estimation (MDE) models are limited by the availability of sufficient and diverse datasets. In the case of MDE models for autonomous driving, this issue is exacerbated by the linearity of the captured data trajectories. We propose a NeRF-based data augmentation pipeline to introduce synthetic data with more diverse viewing directions into training datasets and demonstrate the benefits of our approach to model performance and robustness. Our data augmentation pipeline, which we call NeRFmentation, trains NeRFs on each scene in a dataset, filters out subpar NeRFs based on relevant metrics, and uses them to generate synthetic RGB-D images captured from new viewing directions. In this work, we apply our technique in conjunction with three state-of-the-art MDE architectures on the popular autonomous driving dataset, KITTI, augmenting its training set of the Eigen split. We evaluate the resulting performance gain on the original test set, a separate popular driving dataset, and our own synthetic test set.
Figure 1: NeRFmentation. We propose a novel dataset augmentation pipeline that utilizes NeRF-generated data rendered from poses unseen in the original dataset to improve the robustness of Monocular Depth Estimation networks. The method is intended to improve generalization capability in cases where the initial dataset provides limited spatial variability. The RGB image is taken from the KITTI Dataset [12].
Keywords:
3D from a Single Image Data Augmentation
@@ 1 Introduction
As the field of computer vision continues to advance at a rapid pace, the pursuit of safer and more reliable autonomous driving systems remains paramount. Within this context, monocular depth estimation (MDE) plays an important role, offering a key solution to the complex challenge of depth perception in dynamic road environments. However, it also comes with extra challenges compared to stereo and general multi-camera depth estimation. Accurately estimating the depth from a single image remains a challenging problem due to the inherent scale ambiguity in 2D projections of 3D scenes, which leads to an infinite number of geometrically plausible 3D reprojections for every 2D image. Therefore, monocular depth estimation is an underdetermined problem, which relies on extracting and interpreting visual cues in images correctly. To address this challenge, numerous approaches leverage large-scale datasets to train deep neural networks, allowing them to learn to interpret visual cues and transform them into complex mappings between image features and corresponding depth information.
In order to make monocular depth estimators fit for real-world usage, their generalizability to unseen environments is of utmost importance. However, so far the overwhelming majority of research in this field has been dedicated to improving in-distribution test errors specifically on the KITTI dataset [12]. There is little knowledge available on how state-of-the-art models trained on these datasets perform on zero-shot out-of-distribution evaluation [35, 24, 14].
@@
Furthermore, we believe that improving the generalizability of MDE models does not require the costly and complex creation of new datasets, but can be achieved using dataset augmentation. In recent years, Neural Radiance Field (NeRF) [20] methods have emerged, which can encode complete 3D scenes. NeRFs have the ability to generate photorealistic and geometrically consistent, dense RGB-D images from novel viewpoints with remarkable quality. Therefore, they are powerful enough to use their output to train downstream models, such as monocular depth estimators. Our main contributions are the following:
  • $\bullet$
    We propose NeRFmentation, a simple but effective dataset augmentation pipeline which uses NeRFs to generate high-quality RGB-D images from novel viewpoints to enlarge and diversify existing real-world depth estimation datasets to increase the robustness of trained models.
  • $\bullet$
    We present extensive experiments on how well a variety of state-of-the-art models, trained on the KITTI dataset [12], perform on zero-shot out-of-distribution evaluation, namely on the Waymo Open Dataset [32].
  • $\bullet$
    We show that NeRFmentation outperforms other dataset augmentation techniques both on in- and out-of-distribution test tasks. Furthermore, we provide insights into when classic in-training augmentations (random rotations, flips, and crops) should be used in conjunction with dataset augmentations.
While we showcase the benefit of our data augmentation method for the task of monocular depth estimation, we envision the widespread usage of our approach for a variety of computer vision tasks such as semantic segmentation and surface normals estimation due to its simplicity and effectiveness.
@@ 2 Related work
Monocular Depth Estimation.  This task involves predicting depth values for each pixel in a 2D image. Due to the inherent difficulty in predicting the correct 3D reprojection to a 2D image, well-performing models employ sophisticated data-driven methods to infer the depth values. Various approaches have been developed, including CNN-based encoder-decoder architectures [26, 8] and hybrid models combining CNNs with vision transformers (ViTs) [7]. Some methods treat MDE as a regression problem, while others approach it as an ordinal regression task [9]. Notable examples include AdaBins [5], which uses adaptive bin centers and sizes, DepthFormer [17], which directly encodes images using ViTs, and BinsFormer [18], which combines ordinal regression on binned depth with transformer networks. Our work doesn’t introduce a new MDE architecture but demonstrates the versatility of our method using these existing models [5, 17, 18].
Neural Radiance Fields (NeRFs).  NeRFs are a scene representation method introduced in the seminal work [20] by Mildenhall et al. Using a sparse set of input views, they optimize an underlying volumetric scene function, taking advantage of a differentiable rendering pipeline. Novel views as well as depth data can then be synthesized by querying the network with novel camera poses. Nerfstudio [33] facilitates development of NeRF-based architectures. Nerfacto, an extension of NeRF, incorporates recent advancements like camera pose refinement, per-image appearance conditioning, proposal sampling, scene contraction, and hash encoding. depth-nerfacto adds depth supervision [6], while nerfacto-huge offers the highest fidelity among nerfacto variants.
Zero-shot out-of-distribution model testing.  While zero-shot model testing has been a common practice in the field of natural language processing [41, 25, 34] and other subfields in computer vision [36, 44], only recently has this trend reached monocular depth estimation. Van Dijk et al. [35] analyze the inherent biases towards the training dataset learned by MDEs by introducing synthetic objects into the testing dataset and checking how well the models react to them. MiDaS [24] is one of the works combining different datasets to improve the generalizability of MDEs to unseen data. Yin et al. [40] design a model architecture that allows predicting depth for arbitrary images by projecting different camera setups into a canonical space before estimating the depth and subsequently projecting back. Guizilini et al. [14] train their intrinsics-aware MDE on five datasets and evaluate it on four other datasets and show impressive results. However, they are not able to match current state-of-the-art models trained on these evaluation datasets.
Offline and online data augmentation.  Data augmentation is a well-known regularization technique in Machine Learning that increases generalizability by applying label-preserving transforms to input samples, either during training (online) [29, 15] or before training (offline) [43, 39, 11, 1]. Classical methods for CNNs typically involve simple transformations like random translations, rotations, or flips [29, 15]. Recent advancements in generative modeling, particularly GANs [43, 27, 4, 1] and diffusion models [2], have enabled the offline generation of photo-realistic synthetic data. Various approaches have been developed to improve model performance. Sandfort et al. [27] use CycleGAN for medical CT segmentation, improving performance even for out-of-distribution samples. Atapour-Abarghouei et al. [1] train MDEs on synthetic video game data and evaluate on KITTI using style transfer. Gasperini et al. [10] demonstrate online augmentation for MDE in adverse driving conditions. While recent methods often transfer styles in the image domain [43, 39, 1], our approach uniquely augments the pose space while maintaining image space consistency.
Synthetic datasets and virtual views for MDE.  Rajpal et al. show in [23] that it is possible to improve monocular depth estimation performance on real-world data using a fully synthetic custom dataset. Mancini et al. [19] use two synthetic outdoor datasets to train an MDE and subsequently evaluate it on KITTI, where they achieve results comparable to Eigen et al. [8], as well as their own synthetic datasets. Bauer et al. [3] use in-training novel view synthesis to improve depth prediction. In contrast to [23, 19], we do not use a fully synthetic dataset to improve monocular depth estimators but propose to use an existing dataset and extend it with virtual samples like Bauer et al. [3]. However, our pipeline allows for more diverse poses and can be used independently of the training procedure as we generate samples before the training of the MDE network.
3 Method
Figure 2: Our proposed pipeline: (1) Train NeRFs for each scene in MDE dataset, reserving images for quality evaluation. (2) Filter out subpar NeRFs. (3) Render novel views by perturbing original poses. (4) Combine novel and original views to create NeRFmented training dataset for MDE network. Source dataset: KITTI [12]
We propose a novel dataset augmentation technique for Monocular Depth Estimation which leverages the fact that many real-world image datasets have a sequence-like structure, meaning that the same environment is captured from many slightly different viewpoints. Using the provided pose information, we can build 3D models of the scenes in a dataset, which can subsequently be used to generate new RGB-D pairs from slightly different viewpoints, increasing the size of the original dataset.
3.1 Proposed pipeline
We split each scene in the source dataset into sub-scenes and train a separate depth-supervised NeRF using RGB-D images and their corresponding poses to encode a 3D representation. To ensure a high reconstruction quality for each NeRF, we withhold a small part of the input data from the training. Once the training has concluded, we use the withheld data as a small validation split to potentially filter out scenes with subpar image- or depth reconstruction quality. We construct novel views by applying minor 3D rigid body transformations to the original poses. Then, we use the trained NeRFs to render dense RGB-D images from these novel viewpoints. Finally, we combine the RGB-D images from our novel views with the source dataset to form an augmented dataset and train the downstream monocular depth estimation network on it. The overall pipeline is depicted in Fig. 2.
3.2 Scene representation model
For the NeRFs of our pipeline, we use the off-the-shelf Nerfstudio framework [33], which provides convenient access to various NeRF methods and extensions. Because our method requires high-quality depth reconstruction and image reconstruction, we base our model on a combination of the depth-nerfacto and nerfacto-huge architectures, which were both introduced by Tancik et al. in [33]. This way, we benefit from depth supervision that is included in the depth-nerfacto training while increasing the learning capacity of the network by replacing the MLP of depth-nerfacto with the larger MLP of nerfacto-huge. Fig. 3 shows the qualitative performance gain of our depth-nerfacto-huge compared to the vanilla depth-nerfacto.
@@ 3.2 Scene representation model
Figure 3: Qualitative comparison of depth-nerfacto vs. depth-nerfacto-huge reconstruction on KITTI [12]. The figure shows reconstructions of training images using both depth-nerfacto and depth-nerfacto-huge. On average, depth-nerfacto-huge outputs exhibit higher levels of sharpness and better accuracy in both the generated RGB and depth images compared to depth-nerfacto.
4 Experiments
4.1 Datasets
We consider the following two datasets in our experiments.
KITTI [12] The KITTI dataset is an outdoor dataset designed for diverse autonomous driving scenarios, featuring stereo-image pairs and annotated LiDAR ground-truth depth data. The KITTI Eigen split [8] is a popular dataset split for monocular depth estimation with approximately 23,000 stereo-image pairs for training which are separated into 32 scenes for training and validation. The corresponding validation split consists of 888 stereo-image pairs. The test split contains 697 images from 29 different scenes.
Waymo Open Dataset [32] The Waymo Open Dataset, developed by Waymo LLC for autonomous driving tasks, incorporates data from five high-resolution cameras and five LiDARs. It offers wider camera viewpoints and denser depth maps compared to datasets like KITTI, enabling more robust model training. To evaluate our NeRFmented models’ robustness, we test them on a subset of the Waymo test dataset, using only front, front-left, and front-right facing camera images. This selection minimizes domain shift from standard forward-facing camera MDE datasets like KITTI.
4.2 Dataset augmentation baselines
We evaluate the benefits of using our NeRFmented KITTI dataset for zero-shot MDE on the Waymo test dataset by comparing it with the domain-adaptation methods Fourier Domain Analysis (FDA) [39] and style transfer [1].
For FDA, we only use a single Fourier-domain value of $\beta=0.0025$, which corresponds to replacing the amplitudes of the lowest 0.25 % of the source image frequencies (KITTI) with the amplitudes of the lowest 0.25 % of the target frequencies (Waymo). This way the domain gap is reduced slightly, but no significant artifacts are introduced. Regarding style transfer, we augment the original KITTI Eigen train split with style-transferred RGBs as described in [1], leading to a total dataset size of 29,158 images (+26 %). The ground truth depth maps remain unchanged. Furthermore, we also ablate the effect of using classical data augmentation techniques in the training pipeline consisting of random rotations, flips, and crops.
4.3 Novel view synthesis strategies for data augmentation
We propose and evaluate five data augmentation strategies: (i) Reconstruction: Re-rendering the exact training poses used to train the NeRF. This completely densifies the sparse depth map from the source dataset. (ii) Interpolated: Creating novel poses by interpolating between pairs of training poses. As NeRFs interpolate between embeddings, these renders benefit from good depth and RGB supervision. (iii) Angled: Rendering two additional views for each training pose, rotated horizontally by ±$\pm 3^{\circ}$. This simulates plausible motion performed by a driving car. (iv, v) Translated horizontally, vertically: Generating two additional views for each original pose by translating the camera horizontally/vertically ±$\pm 30\,cm$. For the KITTI Eigen train split [8, 12] this results in dataset sizes of i) 29,836 (+28 %), ii) 29,744 (+28 %), iii) 32,132 (+39 %), iv) 34,949 images (+51 %), and v) 34,949 images (+51 %). An ablation on the optimal amount of NeRFmented data to add is provided in the supplementary.
The motivation behind these data augmentation strategies is to analyze the influence of depth-densification and pose diversification. With this diverse collection, we can evaluate whether it is sufficient to use NeRFs as depth completion networks or whether certain novel views of training scenes can improve performance on evaluation datasets.
GT RGB
Our RGB
GT Depth
Our Depth
Figure 4: Qualitative NeRF reconstruction results on KITTI [12]. Original KITTI images are compared with those generated by trained and filtered NeRFs for matching camera poses. The reconstructed RGB images closely resemble the originals, while NeRFs also complete sparse ground truth depth maps.
4.4 Training of NeRFs
We train a depth-nerfacto-huge model for each scene of the KITTI Eigen [12, 8] train split and retain 10 % of the training data as a validation split. As we experimentally find that the reconstruction quality of NeRFs drastically reduces as the size of a scene increases, the varying sizes of the scenes in the KITTI dataset require splitting them into sub-scenes. However, it is unfavorable to decrease the scene size excessively, as this would increase the computational complexity without a gain in quality. To achieve a good balance, we ensure that no pose in each sub-scene is further than 50 m away from the first pose in that sub-scene. For each sub-scene, we train a separate NeRF.
For each NeRF, we tune the weight of the depth loss and whether the poses are refined or not by running a grid search for 10,000 steps. We observe that when using NeRFs for 3D reconstruction, there is a strong trade-off between RGB quality and accuracy of the geometry, which in turn affects the accuracy of the depth maps. Since we require both to be of high quality, we use the small validation split that we withheld from the training and re-render the RGB-D images for each of their poses using all NeRFs from our grid search. Then, we select the NeRF that minimizes a trade-off validation measure $T_{\text{RGBD}}=\sqrt{\alpha\cdot\text{Abs. Rel}^{2}+\mathrm{LPIPS}^{2}}$, where we use the absolute relative error of the reprojected depth maps as a quality measure of the geometry and the LPIPS score [42] as a quality measure of the RGB reconstruction. During our experiments, we set $\alpha=10$.
Finally, we continue training the best run of the hyperparameter tuning of each NeRF until we reach 30,000 iterations. Then, we apply a filtering step that eliminates NeRFs that fail to meet a specified dataset-dependent reconstruction quality standard. To achieve this, we utilize the trained NeRF for each scene to render the RGB-D images from the poses of the small withheld validation split and compare them to the corresponding ground-truth RGB-D images. This involves computing the absolute relative error to assess the depth reconstruction quality and the LPIPS score [42] to evaluate the image reconstruction quality. We discard all NeRFs that do not achieve an LPIPS score of less than $0.22$ or an absolute relative error of less than $0.05$ on the 3D reconstructions. The threshold for LPIPS is chosen empirically by observing which value produces minimally visually acceptable RGB images. The threshold for depth absolute relative error is chosen because it is an approximate lower bound of state-of-the-art MDE networks. We filter any scenes with a higher score so as not to introduce extraneous noise into the training dataset.
Training a depth-nerfacto-huge model on a sub-scene takes approximately 45 min on an RTX 4090. In our case, we obtained 434 sub-scenes from KITTI. Therefore, our total GPU compute for training the NeRFs was around 2 weeks. Generating a novel view at the native KITTI resolution takes around 1.4 seconds per frame. We note that both training the NeRFs and generating novel views can easily be parallelized on a sub-scene level when multiple GPUs are available. As this is a data augmentation strategy, there is no performance penalty when performing inference.
Table 1: Comparison of data augmentation methods for models trained on augmented KITTI datasets evaluated on the Waymo dataset. We showcase the performance of baseline MDEs and compare them to the NeRFmented (Ours) versions of the same models evaluated against the unseen Waymo dataset. For NeRFmentation, the NeRFm. Vert. ±$\pm 30\,cm$ strategy is shown. CAug indicates whether classic augmentations were used during training. The best results are in bold. The second best results are underlined, and the baseline is highlighted.